# Solving the Poisson equation using coupled Markov chains

### v3

@@ B.4 Unbiased approximation of $\pi(h)$
To determine that $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]<\infty$ for $1\geq\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$ we define $S_{1}=\omega_{K}\xi_{K}^{-1}h(Z_{K})$ where $K\mid\hat{\pi}\sim{\rm Categorical}\{\xi_{1},\ldots,\xi_{N}\}$. Then $S_{R}$ is less than $S_{1}$ in the convex order, i.e. $S_{R}\leq_{{\rm cx}}S_{1}$ (see, e.g., Shaked & Shanthikumar 2007) for any $R\geq 2$ since one may define $S_{1}$ by drawing $K\sim\frac{1}{R}\sum_{i=1}^{R}\delta_{K_{i}}$ and it is then clear that $\mathbb{E}[S_{1}\mid\sigma(K_{1:R},\omega_{1:N},Z_{1:N})]=S_{R}$. It follows that $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]\leq\mathbb{E}\left[\left|S_{1}\right|^{p}\right]$ since $x\mapsto\left|x\right|^{p}$ is convex for $p\geq 1$. By Lemma 44, $\hat{\pi}$ is defined so that
$\max_{i\in\{1,\ldots,N\}}\left|\omega_{i}\right|\leq\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right),$
and $N\leq\ell-k+1+2\tau$. We have by Hölder’s inequality
$\mathbb{E}\left[\left|S_{1}\right|^{p}\right]$ $=\mathbb{E}\left[\left|\omega_{K}\xi_{K}^{-1}h(Z_{K})\right|^{p}\right]$
$\leq a^{-p}\mathbb{E}\left[\left|N\omega_{K}h(Z_{K})\right|^{p}\right]$
$\leq a^{-p}\left\{\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right)\right\}^{p}\mathbb{E}\left[\left|N\right|^{\kappa}\right]^{\frac{p}{\kappa}}\mathbb{E}\left[\left|h(Z_{K})\right|^{\frac{\kappa p}{\kappa-p}}\right]^{1-\frac{p}{\kappa}},$
and since $\mathbb{E}\left[\left|N\right|^{\kappa}\right]<\infty$ by Assumption 2 it remains to show that $\mathbb{E}\left[\left|h(Z_{K})\right|^{\frac{\kappa p}{\kappa-p}}\right]<\infty$. Now, since $\mathbb{P}(K=i\mid\sigma(N,Z_{1},\ldots,Z_{N}))\leq bi^{-1}$ for all $i\geq 1$, and taking any $\varepsilon\in(0,\kappa)$, we obtain for $q=\kappa p/(\kappa-p)$,
$\mathbb{E}\left[\left|h(Z_{K})\right|^{q}\right]$ $=\sum_{i=1}^{\infty}\mathbb{E}\left[\mathbb{I}(K=i)\left|h(Z_{i})\right|^{q}\right]$
$=\sum_{i=1}^{\infty}\mathbb{E}\left[\mathbb{I}(N\geq i,K=i)\left|h(Z_{i})\right|^{q}\right]$
$\leq b\sum_{i=1}^{\infty}\frac{1}{i}\mathbb{E}\left[\mathbb{I}(N\geq i)\left|h(Z_{i})\right|^{q}\right]$
@@
For $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, it follows that we may take $\varepsilon=\kappa(m-q)/m$ so that $\kappa q/(\kappa-\varepsilon)=m$. Moreover, since $Z_{i}\sim\pi_{0}P^{t}$ for some $t\in\mathbb{N}$, $\mathbb{E}\left[\left|h(Z_{i})\right|^{m}\right]\leq M\pi(\left|h\right|^{m})$ by Lemma 35 and so using the fact that $(1-\frac{\varepsilon}{\kappa})(1-\frac{p}{\kappa})=\frac{p}{m}$, we obtain
$\mathbb{E}\left[\left|S_{1}\right|^{p}\right]$ $\leq a^{-p}\left\{\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right)\right\}^{p}\mathbb{E}\left[\left|N\right|^{\kappa}\right]^{\frac{p}{\kappa}}\left\{b\mathbb{E}\left[N^{\kappa}\right]^{\frac{\varepsilon}{\kappa}}M^{1-\frac{\varepsilon}{\kappa}}\pi(\left|h\right|^{m})^{1-\frac{\varepsilon}{\kappa}}\sum_{i=1}^{\infty}\frac{1}{i^{1+\varepsilon}}\right\}^{1-\frac{p}{\kappa}}$
@@
which is finite and from which we may conclude. ∎
Example 46.
Natural choices of $\xi_{i}$ are to take $\xi_{i}=1/N$ or $\xi_{i}\propto\left|\omega_{i}\right|$. In the latter case, it follows from Lemma 44 that
@@
and so one may take $a=1/(1+\frac{\ell-k}{L})$ and $b=1+(\ell-k)/L$ in Proposition 45 for this choice.
@@ B.5 Unbiased approximation of $\pi(h_{0}\cdot g)$
We now look at combinations of unbiased fishy function estimation and unbiased estimation of $\pi(h)$. This will involve estimators $G_{y}(x)$ at several points $x$ drawn randomly. We first define an alternative representation that will avoid some ambiguity in the following developments. We define a probability measure $Q$ such that, with $U\sim Q$,
@@
and we define an extended distribution $\check{\pi}({\rm d}x,{\rm d}u)=\pi({\rm d}x)Q({\rm d}u)$ with $\check{\pi}$-invariant Markov kernel $T(x,u;{\rm d}y,{\rm d}v)=P(x,{\rm d}y)Q({\rm d}v)$, and coupled Markov kernel
$\bar{T}(x,u,y,v;{\rm d}x^{\prime},{\rm d}u^{\prime},{\rm d}y^{\prime},{\rm d}v^{\prime})=\bar{P}(x,y;{\rm d}x^{\prime},{\rm d}y^{\prime})Q({\rm d}u^{\prime})\cdot\left\{\mathbb{I}(x^{\prime}=y^{\prime})\delta_{u^{\prime}}({\rm d}v^{\prime})+\mathbb{I}(x^{\prime}\neq y^{\prime})Q({\rm d}v^{\prime})\right\}.$
An interpretation of this Markov kernel is that if $(X^{\prime},U^{\prime},Y^{\prime},V^{\prime})\sim\bar{T}(x,y,u,v;\cdot)$ then $(X^{\prime},Y^{\prime})\sim\bar{P}(x,y;\cdot)$ and if $X^{\prime}=Y^{\prime}$ then $V^{\prime}=U^{\prime}\sim Q$ but if $X^{\prime}\neq Y^{\prime}$ then $U^{\prime},V^{\prime}\sim Q$ independently.
We use $\check{G}^{f}$ to indicate that $\check{G}_{y,v}^{f}(x,u)$ is an unbiased approximation of the fishy function for $T$ and $f$ as opposed to $P$ and $h$, i.e. such that
@@
where $\check{g}_{\star}^{f}(x,u):=\sum_{t=0}^{\infty}T^{t}f_{0}(x,u)$, with $f_{0}:=f-\check{\pi}(f)$, is the fishy function for $T$ and $f$ in $L_{0}^{1}(\check{\pi})$.
The Markov chain $(X,U,Y,V)$ has the same meeting time as the Markov chain $(X,Y)$ by construction, so Assumption 2 holds for this chain with $\pi$ replaced by $\check{\pi}$. Similarly if ${\rm d}\mu/{\rm d}\pi\leq M$ then with $\check{\mu}=\mu\otimes Q$ and $\check{\pi}=\pi\otimes Q$ we have ${\rm d}\check{\mu}/{\rm d}\check{\pi}\leq M$. Hence, we may apply Proposition 43 or Proposition 45 to deduce lack-of-bias of an appropriate approximation of $\check{\pi}(\phi)$ and finite $p$th moments if $\phi\in L^{m}(\check{\pi})$ and $1\geq\frac{1}{p}>\frac{1}{\kappa}+\frac{1}{m}$.
The following two lemmas provide conditions for $g_{y}\in L^{q}(\check{\pi})$ and $h\cdot g_{y}\in L^{s}(\check{\pi})$, which are used to analyze both the MCMC estimator of $v(P,h)$ of Section 4.2 and the unbiased estimators of Section 4.3.
Lemma 47.
Let $h\in L^{m}(\pi)$ for some $m>1$. Under Assumption 2, for $\pi$-almost all $y$, $g_{y}\in L^{q}(\check{\pi})$ for $q\geq 1$ such that $\frac{1}{q}>\frac{1}{m}+\frac{1}{\kappa}$.
@@
Taking $\gamma=\pi\otimes\delta_{y}$ in Theorem 37 we obtain that, for $\pi$-almost all $y$,
$\check{\pi}(\left|g\right|^{q})^{\frac{1}{q}}=\mathbb{E}_{\pi}\left[\left|G_{y}(X_{0})\right|^{q}\right]^{\frac{1}{q}}\leq\zeta\left(\frac{(m-q)\kappa}{mq}\right)\mathbb{E}_{\pi,y}[\tau^{\kappa}]^{\frac{m-q}{mq}}\left\{\left\|h\right\|_{L^{m}(\pi)}+\psi(h,m,\delta_{y})\right\}<\infty.$
Lemma 48.
Let $h\in L^{m}(\pi)$ for some $m>1$. Under Assumption 2, with $\phi_{1}(x,u)=h(x)g_{y}(x,u)$, for $\pi$-almost all $y$, $\phi_{1}\in L^{s}(\check{\pi})$ for $s\geq 1$ such that $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$.
@@
$\check{\pi}(\left|\phi_{1}\right|^{s})$ $\leq\check{\pi}(\left|f\right|^{\frac{s}{\delta}})^{\delta}\check{\pi}(\left|g\right|^{\frac{s}{1-\delta}})^{1-\delta}=\pi(\left|h\right|^{\frac{s}{\delta}})^{\delta}\check{\pi}(\left|g\right|^{\frac{s}{1-\delta}})^{1-\delta}.$
With $q=s/(1-\delta)$, we deduce by Lemma 47 that if $\frac{1}{q}>\frac{1}{m}+\frac{1}{\kappa}$ then $\check{\pi}(\left|g_{y}\right|^{q})<\infty$ for $\pi$-almost all $y$. On the other hand, if $s<m\delta$, then $\pi(\left|h\right|^{\frac{s}{\delta}})<\infty$. Taking $\delta=\kappa/(2\kappa+m)$ we find that $m\delta=(1-\delta)(\frac{1}{m}+\frac{1}{\kappa})^{-1}$, and this implies that $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$ is sufficient for $\check{\pi}(\left|\phi_{1}\right|^{s})<\infty$. ∎
Definition 49 (Estimator of $\pi(h_{0}\cdot g_{y})$).
Let $H$ be an unbiased estimator of $\pi(h)$ as defined in Proposition 43, with ${\rm d}\pi_{0}/{\rm d}\pi\leq M$. Let $\phi_{H}(x,u)=(h(x)-H)g_{y}(x,u)$ and, with random variables independent to those used to define $H$, let $\Phi_{R}$ (resp. $\Phi_{0}$) be the approximation corresponding to $S_{R}$ (resp. $H_{k:\ell}$) of $\check{\pi}(\phi_{H})$ in Proposition 45 (resp. Proposition 43) with the unbiased signed measure approximating $\check{\pi}$ involving random variables independent to those used to define $H$.
Proposition 50.
Under Assumption 2 with $\kappa>2$, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-2)$. Then for any $R\in\{0,1,\ldots\}$,
@@
We first determine $s\geq 1$ such that $\phi_{H}\in L^{s}(\check{\pi})$. Let $\phi_{1}(x,u)=h(x)g_{y}(x,u)$ and $\phi_{2,c}(x,u)=cg_{y}(x,u)$, so that $\phi_{H}=\phi_{1}-\phi_{2,H}$. We find that for a fixed $H$ and $s\geq 1$,
$\check{\pi}(\left|\phi_{H}\right|^{s})^{\frac{1}{s}}\leq\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{1}{s}}+\check{\pi}(\left|\phi_{2,H}\right|^{s})^{\frac{1}{s}}=\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{1}{s}}+\left|H\right|\check{\pi}(\left|g_{y}\right|^{s})^{\frac{1}{s}}.$
By Lemma 47, $\check{\pi}(\left|g_{y}\right|^{s})<\infty$ if $\frac{1}{s}>\frac{1}{m}+\frac{1}{\kappa}$. By Lemma 48, $\check{\pi}(\left|\phi_{1}\right|^{s})<\infty$ if $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$. Hence, $\phi_{H}\in L^{s}(\check{\pi})$ for $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$. It follows that if $\frac{1}{p}>\frac{2}{m}+\frac{2}{\kappa}$ then there exists $s\in[1,(\frac{2}{m}+\frac{1}{\kappa})^{-1})$ such that by Proposition 43,
@@
$\mathbb{E}\left[\left\|\phi_{H}\right\|_{L^{s}(\check{\pi})}^{p}\right]$ $\leq\mathbb{E}\left[\left\{\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{1}{s}}+\left|H\right|\check{\pi}(\left|g_{y}\right|^{s})^{\frac{1}{s}}\right\}^{p}\right]$
$\leq 2^{p}\left\{\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{p}{s}}+\mathbb{E}\left[\left|H\right|^{p}\right]\check{\pi}(\left|g_{y}\right|^{s})^{\frac{p}{s}}\right\},$
and $\mathbb{E}\left[\left|H\right|^{p}\right]<\infty$ by Proposition 43 if $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$ , which imposes no additional constraints on $p$ or $s$. Hence, we may conclude that for $\frac{1}{p}>\frac{2}{m}+\frac{2}{\kappa}$, $\mathbb{E}\left[\left|\Phi_{R}\right|^{p}\right]<\infty$ for $R\geq 0$. For the lack-of-bias property, we consider $p=1$ and if $m>2\kappa/(\kappa-2)$ then, by Proposition 43 and Proposition 45,

### v4

@@ B.4 Unbiased approximation of $\pi(h)$
To determine that $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]<\infty$ for $1\geq\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$ we define $S_{1}=\omega_{I}\xi_{I}^{-1}h(Z_{I})$ where $I\mid\hat{\pi}\sim{\rm Categorical}\{\xi_{1},\ldots,\xi_{N}\}$. Then $S_{R}$ is less than $S_{1}$ in the convex order, i.e. $S_{R}\leq_{{\rm cx}}S_{1}$ (see, e.g., Shaked & Shanthikumar 2007) for any $R\geq 2$ since one may define $S_{1}$ by drawing $I\sim{R}^{-1}\sum_{r=1}^{R}\delta_{I_{r}}$ and it is then clear that $\mathbb{E}[S_{1}\mid\sigma(I_{1:R},\omega_{1:N},Z_{1:N})]=S_{R}$. It follows that $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]\leq\mathbb{E}\left[\left|S_{1}\right|^{p}\right]$ since $x\mapsto\left|x\right|^{p}$ is convex for $p\geq 1$. Using Lemma B.21, the fact that $N\leq\ell-k+1+2\tau$ and Hölder’s inequality,
$\mathbb{E}\left[\left|S_{1}\right|^{p}\right]$ $=\mathbb{E}\left[\left|\omega_{I}\xi_{I}^{-1}h(Z_{I})\right|^{p}\right]$
$\leq a^{-p}\mathbb{E}\left[\left|N\omega_{I}h(Z_{I})\right|^{p}\right]$
$\leq a^{-p}\left\{\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right)\right\}^{p}\mathbb{E}\left[\left|N\right|^{\kappa}\right]^{\frac{p}{\kappa}}\mathbb{E}\left[\left|h(Z_{I})\right|^{\frac{\kappa p}{\kappa-p}}\right]^{1-\frac{p}{\kappa}},$
and since $\mathbb{E}\left[\left|N\right|^{\kappa}\right]<\infty$ by Assumption 2 it remains to show that $\mathbb{E}\left[\left|h(Z_{I})\right|^{\frac{\kappa p}{\kappa-p}}\right]<\infty$. Now, since $\mathbb{P}(I=i\mid\sigma(N,Z_{1},\ldots,Z_{N}))\leq bi^{-1}$ for all $1\leq i\leq N$, and taking any $\varepsilon\in(0,\kappa)$, we obtain for $q=\kappa p/(\kappa-p)$,
$\mathbb{E}\left[\left|h(Z_{I})\right|^{q}\right]$ $=\sum_{i=1}^{\infty}\mathbb{E}\left[\mathds{1}(N\geq i,I=i)\left|h(Z_{i})\right|^{q}\right]$
$\leq b\sum_{i=1}^{\infty}\frac{1}{i}\mathbb{E}\left[\mathds{1}(N\geq i)\left|h(Z_{i})\right|^{q}\right]$
@@
For $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, it follows that we may take $\varepsilon=\kappa(m-q)/m$ so that $\kappa q/(\kappa-\varepsilon)=m$. Moreover, since $Z_{i}\sim\pi_{0}P^{t}$ for some $t\in\mathbb{N}$, $\mathbb{E}\left[\left|h(Z_{i})\right|^{m}\right]\leq M\pi(\left|h\right|^{m})$ by Lemma B.12 and so using the fact that $(1-\frac{\varepsilon}{\kappa})(1-\frac{p}{\kappa})=\frac{p}{m}$, we obtain
$\mathbb{E}\left[\left|S_{1}\right|^{p}\right]$
$\leq a^{-p}\left\{\frac{1}{\ell-k+1}\left(1+\frac{\ell-k}{L}\right)\right\}^{p}\mathbb{E}\left[\left|N\right|^{\kappa}\right]^{\frac{p}{\kappa}}\left\{b\mathbb{E}\left[N^{\kappa}\right]^{\frac{\varepsilon}{\kappa}}M^{1-\frac{\varepsilon}{\kappa}}\pi(\left|h\right|^{m})^{1-\frac{\varepsilon}{\kappa}}\sum_{i=1}^{\infty}\frac{1}{i^{1+\varepsilon}}\right\}^{1-\frac{p}{\kappa}}$
@@
which is finite. We thus conclude. ∎
Example B.23.
Natural choices of $\xi_{i}$ are to take $\xi_{i}=1/N$ or $\xi_{i}\propto\left|\omega_{i}\right|$. In the latter case, it follows from Lemma B.21 that
@@
and so one may take $a=1/(1+\frac{\ell-k}{L})$ and $b=1+(\ell-k)/L$ in Proposition B.22 for this choice.
@@ B.5 Unbiased approximation of $\pi(h_{0}\cdot g)$
We now look at combinations of unbiased fishy function estimation and unbiased estimation of $\pi(h)$. This will involve estimators $G_{y}(x)$ at random points $x$. We first introduce an alternative representation to avoid ambiguity in the following developments. We define a probability measure $Q$ such that, with $U\sim Q$,
@@
with $G_{y}(x)$ in Definition 5, and we define an extended distribution $\check{\pi}({\rm d}x,{\rm d}u)=\pi({\rm d}x)Q({\rm d}u)$ with $\check{\pi}$-invariant Markov kernel $T(x,u;{\rm d}y,{\rm d}v)=P(x,{\rm d}y)Q({\rm d}v)$, and coupled Markov kernel
$\bar{T}(x,u,y,v;{\rm d}x^{\prime},{\rm d}u^{\prime},{\rm d}y^{\prime},{\rm d}v^{\prime})=$ $\bar{P}(x,y;{\rm d}x^{\prime},{\rm d}y^{\prime})\,Q({\rm d}u^{\prime})$
×$\times\left\{\mathds{1}(x^{\prime}=y^{\prime})\,\delta_{u^{\prime}}({\rm d}v^{\prime})+\mathds{1}(x^{\prime}\neq y^{\prime})\,Q({\rm d}v^{\prime})\right\}.$
Thus, if $(X^{\prime},U^{\prime},Y^{\prime},V^{\prime})\sim\bar{T}(x,y,u,v;\cdot)$ then $(X^{\prime},Y^{\prime})\sim\bar{P}(x,y;\cdot)$ and if $X^{\prime}=Y^{\prime}$ then $V^{\prime}=U^{\prime}\sim Q$ but if $X^{\prime}\neq Y^{\prime}$ then $U^{\prime},V^{\prime}\sim Q$ independently.
We denote by $\check{G}_{y,v}^{f}(x,u)$ an unbiased approximation of a fishy function associated with the transition $T$ and test function $f$, as opposed to $P$ and $h$, i.e.
@@
where $\check{g}_{\star}^{f}(x,u):=\sum_{t=0}^{\infty}T^{t}f_{0}(x,u)$, with $f_{0}:=f-\check{\pi}(f)$ for $f$ in $L_{0}^{1}(\check{\pi})$.
The Markov chain $(X,U,Y,V)$ has the same meeting time as the Markov chain $(X,Y)$ by construction, so Assumption 2 holds for this chain with $\pi$ replaced by $\check{\pi}$. Similarly if ${\rm d}\mu/{\rm d}\pi\leq M$ then with $\check{\mu}=\mu\otimes Q$ and $\check{\pi}=\pi\otimes Q$ we have ${\rm d}\check{\mu}/{\rm d}\check{\pi}\leq M$. Hence, we may apply Proposition B.20 or Proposition B.22 to deduce lack-of-bias of an appropriate approximation of $\check{\pi}(\phi)$ and finite $p$th moments if $\phi\in L^{m}(\check{\pi})$ and $1\geq\frac{1}{p}>\frac{1}{\kappa}+\frac{1}{m}$.
The following two lemmas provide conditions for $\bar{g}_{y}\in L^{q}(\check{\pi})$ and $h\cdot\bar{g}_{y}:(x,u)\mapsto h(x)\bar{g}_{y}(x,u)\in L^{s}(\check{\pi})$, which are used to analyze both the MCMC estimator of $v(P,h)$ of Section 4.2 (EPAVE) and the unbiased estimators of Section 4.3 (UPAVE).
Lemma B.24.
Let $h\in L^{m}(\pi)$ for some $m>1$. Under Assumption 2, for $\pi$-almost all $y$, $\bar{g}_{y}\in L^{q}(\check{\pi})$ for $q\geq 1$ such that $\frac{1}{q}>\frac{1}{m}+\frac{1}{\kappa}$.
@@
We observe that $\check{\pi}(\left|\bar{g}_{y}\right|^{q})^{\frac{1}{q}}=\mathbb{E}_{\pi}\left[\left|G_{y}(X_{0})\right|^{q}\right]^{\frac{1}{q}}$ and then we apply Theorem B.14, part 3., with $\gamma=\pi\otimes\delta_{y}$ for $\pi$-almost all $y$. ∎
Lemma B.25.
Let $h\in L^{m}(\pi)$ for some $m>1$. Under Assumption 2, with $\phi_{1}(x,u)=h(x)\bar{g}_{y}(x,u)$, for $\pi$-almost all $y$, $\phi_{1}\in L^{s}(\check{\pi})$ for $s\geq 1$ such that $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$.
@@
$\check{\pi}(\left|\phi_{1}\right|^{s})$ $\leq\check{\pi}(\left|f\right|^{\frac{s}{\delta}})^{\delta}\check{\pi}(\left|\bar{g}_{y}\right|^{\frac{s}{1-\delta}})^{1-\delta}=\pi(\left|h\right|^{\frac{s}{\delta}})^{\delta}\check{\pi}(\left|\bar{g}_{y}\right|^{\frac{s}{1-\delta}})^{1-\delta}.$
With $q=s/(1-\delta)$, we deduce by Lemma B.24 that if $\frac{1}{q}>\frac{1}{m}+\frac{1}{\kappa}$ then $\check{\pi}(\left|\bar{g}_{y}\right|^{q})<\infty$ for $\pi$-almost all $y$. On the other hand, if $s<m\delta$, then $\pi(\left|h\right|^{\frac{s}{\delta}})<\infty$. Taking $\delta=\kappa/(2\kappa+m)$ we find that $m\delta=(1-\delta)(\frac{1}{m}+\frac{1}{\kappa})^{-1}$, and this implies that $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$ is sufficient for $\check{\pi}(\left|\phi_{1}\right|^{s})<\infty$. ∎
Definition B.26 (Estimator of $\pi(h_{0}\cdot g_{y})$).
Let $H$ be an unbiased estimator of $\pi(h)$ as defined in Definition 16. Let $\phi_{H}(x,u)=(h(x)-H)\bar{g}_{y}(x,u)$ and, with random variables independent to those used to define $H$, let $\Phi_{R}$ (resp. $\Phi_{0}$) be the approximation corresponding to $S_{R}$ (resp. $H_{k:\ell}$) of $\check{\pi}(\phi_{H})$ in Proposition B.22 (resp. Proposition B.20) with the unbiased signed measure approximating $\check{\pi}$ involving random variables independent to those used to define $H$.
Proposition B.27.
Under Assumption 2 with $\kappa>2$, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-2)$, and ${\rm d}\pi_{0}/{\rm d}\pi\leq M$. For any $R\in\{0,1,\ldots\}$, $\Phi_{R}$ in Definition B.26 satisfies,
@@
We first determine $s\geq 1$ such that $\phi_{H}\in L^{s}(\check{\pi})$. Let $\phi_{1}(x,u)=h(x)\bar{g}_{y}(x,u)$ and $\phi_{2,c}(x,u)=c\bar{g}_{y}(x,u)$, so that $\phi_{H}=\phi_{1}-\phi_{2,H}$. We find that for a fixed $H$ and $s\geq 1$,
$\check{\pi}(\left|\phi_{H}\right|^{s})^{\frac{1}{s}}\leq\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{1}{s}}+\check{\pi}(\left|\phi_{2,H}\right|^{s})^{\frac{1}{s}}=\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{1}{s}}+\left|H\right|\check{\pi}(\left|\bar{g}_{y}\right|^{s})^{\frac{1}{s}}.$
By Lemma B.24, $\check{\pi}(\left|\bar{g}_{y}\right|^{s})<\infty$ if $\frac{1}{s}>\frac{1}{m}+\frac{1}{\kappa}$. By Lemma B.25, $\check{\pi}(\left|\phi_{1}\right|^{s})<\infty$ if $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$. Hence, $\phi_{H}\in L^{s}(\check{\pi})$ for $\frac{1}{s}>\frac{2}{m}+\frac{1}{\kappa}$. It follows that if $\frac{1}{p}>\frac{2}{m}+\frac{2}{\kappa}$ then there exists $s\in[1,(\frac{2}{m}+\frac{1}{\kappa})^{-1})$ such that by Proposition B.20,
@@
$\mathbb{E}\left[\left\|\phi_{H}\right\|_{L^{s}(\check{\pi})}^{p}\right]$ $\leq\mathbb{E}\left[\left\{\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{1}{s}}+\left|H\right|\check{\pi}(\left|\bar{g}_{y}\right|^{s})^{\frac{1}{s}}\right\}^{p}\right]$
$\leq 2^{p-1}\left\{\check{\pi}(\left|\phi_{1}\right|^{s})^{\frac{p}{s}}+\mathbb{E}\left[\left|H\right|^{p}\right]\check{\pi}(\left|\bar{g}_{y}\right|^{s})^{\frac{p}{s}}\right\},$
and $\mathbb{E}\left[\left|H\right|^{p}\right]<\infty$ by Proposition B.20 if $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$ , which imposes no additional constraints on $p$ or $s$. Hence, we may conclude that for $\frac{1}{p}>\frac{2}{m}+\frac{2}{\kappa}$, $\mathbb{E}\left[\left|\Phi_{R}\right|^{p}\right]<\infty$ for $R\geq 0$. For the lack-of-bias property, we consider $p=1$ and if $m>2\kappa/(\kappa-2)$ then, by Proposition B.20 and Proposition B.22,